0/1pts Incorrect Question 13 How many isomers (constitutional and stereoisomers) exist for dimethylcyclohexane? 1/1 pts Question 14
0/1pts Incorrect Question 16 Which of the following structures represent the same stereoisomer? Br Br CHs H. CH H CH HO CHS Br HO CH 1 3 2 only 1 and 3 only 2 and 3 only I and 2 1, 2 and 3
Answer #1

сHg -сHg 12-рylegalo heтена pos :рos Сраn enantiome Сetg) Cmuo) prtical Tatal no of isomers = 2+1=3 c 2 ко -of а.ror o-1 o(2сHg rens б-й rtimen ala cis СhiaenHiealt Ttal no.of itomers z 2+ l=3 С3 2 па2 ко -f в.f ar O-I 3t2*-1) *(22-1) 4-13Also cos H K pos. Imes meso) mero) iduntient iatica Tetal no of isomers = 2 C 2 In-1) n2,-An yeelohıxane Нens, it ha he ehercoisos.

So dimethylcyclohexane have 9 isomers

We have to submit only one question so plz submit different questions separately .....hv a nice day